A person has invested shs 5000 in the stock market. At the end of the first year the amount has grown to shs 6250, he has had a 25 percent profit. If at the end of the second year his principal has grown to shs 8750, the rate of increase is 40 percent for the year. What is the average rate of increase of his investment during the two years.


1
Expert's answer
2021-07-02T11:39:54-0400
"8750=5000(1+\frac{p}{100})^2"

"(1+\frac{p}{100})^2=1.75"

"p=32.3"